OpenStudy (anonymous):

solve the system of the equation what is the solution? 7u+v=27 7u=+43

OpenStudy (luffingsails):

Take the second equation and solve it for u. Then substitute the result into the first equation and solve it for v. u = 43/7 (first part of solution) Substitute into first equation 7(43/7) + v = 27 43 +v = 27 v = -16 So, u = 43/7, v = -16

OpenStudy (anonymous):

im lost it the answer would then be -16 right for v but what about u

OpenStudy (luffingsails):

The answer for u is given solely by the second equation. If I gave you: 7u = 43 What would u be?

OpenStudy (anonymous):

6.14

OpenStudy (anonymous):

right ????

OpenStudy (luffingsails):

Well, yes, if you are required to provide an exact answer. Otherwise, the answer is 43/7.

OpenStudy (luffingsails):

But after I know what u is, I can substitute that value into equation 1, 7u+v=27. Then I have a single unknown and can solve!